# DocLayLLM: An Efficient and Effective Multi-modal Extension of Large Language Models for Text-rich Document Understanding

### v2

@@ 3.2 CoT Pre-training
During the pre-training phase, we incorporate a wide range of tasks (as detailed in Table 1) to enhance the LLM’s understanding of the visual and layout information inspired by (Luo et al., 2024). Since the datasets related to these tasks are not designed explicitly for instruction-tuning, their annotations are not in a QA format. Therefore, we manually create instructions based on these annotations for training. However, simply converting task descriptions into $Question$ and annotations into $Answer$ is highly inefficient regarding data utilization. To address this, we incorporate a CoT approach during the conversion to instruction-tuning data for specific tasks, asking LLM to solve our problem step-by-step, enhancing the model’s comprehension abilities. Our designed CoT primarily falls into the following three categories:
@@ 3.2.1 Layout-Type-Related CoT
We conduct a layout-type-related CoT for the layout analysis task. To determine the layout type of a given area, it is beneficial first to locate it within the document and then infer based on the nearby text lines and the overall document layout. When giving an area bounded by a bounding box, our layout-type-related CoT can be formulated as following steps:
@@
    Identify the text within the area and, based on this, determine its approximate location within the document (e.g., top left, top right part of the document, etc).
@@ 3.2.2 Table-Structure-Aware CoT
Table-structure-aware CoT is primarily employed in the table analysis task. We have integrated the XY-Cut algorithm Nagy et al. (1992)—a widely-used technique for analyzing table structures. For instance, to locate an element at the $i$-th row and $j$-th column, the CoT process is outlined as follows:
@@
    Identify all the table heads and their respective locations. The head of the $j$-th column corresponds to the $j$-th element in the heads list.
@@
    Output all elements in the $j$-th column when told that the elements of the same column are generally vertically aligned with their corresponding heads.
@@
Applying the above CoT, we break down a complex question into simple steps using the basic XY-Cut thought. This approach improves the LLM’s inference ability through this breakdown and enhances data utilization by effectively leveraging the original annotations.
@@ 3.2.3 Geometry-Related CoT
This is used in geometric analysis, targeting the identification of geometric relationships between two text lines. It improves LLM’s understanding of layout features and enhances inference with geometric relations. Specifically, when giving the text content of two lines:
@@
    Recover the bounding box of the given text lines.
@@
    Calculate the center point of each text line and determine their relative location by comparing their center points.
@@
This design enhances the LLM’s arithmetic reasoning capabilities and helps it more accurately grasp the spatial relationships between different boxes, and also contributes to layout-type-aware CoT.
@@ 3.3 CoT Annealing
Data annealing involves gradually increasing the proportion of high-quality data towards the end of supervised fine-tuning to improve data utilization efficiency and enhance the model’s reasoning abilities. This technique has been proven highly effective and is widely adopted in models like MiniCPM Hu et al. (2024b), Data-CompLM (Li et al., 2024a), and Llama-3.1 (Llama Team, 2024).
During the SFT stage, we employ data with a layout-aware CoT Luo et al. (2024) that helps focus on potential regions of questions and answers within documents, thereby better leveraging the layout features learned during pre-training. It also allows the model to enhance its reasoning ability through step-by-step thinking. However, in the latter stages of SFT, our goal for LLM is to output precise and direct answers to questions. At this point, using CoT data may introduce unnecessary noise with a long response, and the overly uniform layout-aware CoT data might partially compromise the LLM’s generalization abilities.
To balance the strengths and weaknesses of such data, we propose CoT annealing, which reconsiders this data annealing issue from a CoT perspective. Specifically, we generate direct-answer data without CoT from the CoT data, treating it as high-quality data. During the SFT process, we gradually increase the proportion of this direct-answer data, thereby enhancing the LLM’s ability to respond with a clear and direct answer. We initially utilize only the data with CoT, gradually adjusting the ratio of CoT to non-CoT data as the SFT process progresses. Ultimately, we conclude the training using only the data without CoT.
Method Document-oriented VQA KIE
DocVQA VisualMRC Avg. FUNSD- CORD- SROIE- Avg.
Text
Llama2-7B-Chat (Touvron et al., 2023) 64.99 52.84 58.92 48.20 47.70 68.97 54.96
Llama3-8B-Instruct (Llama Team, 2024) 51.79 47.77 49.78 68.57 52.31 61.24 60.70
Text + Layout
Llama2-7B-Chatlayout (Touvron et al., 2023) 56.55 49.26 52.91 58.34 50.93 51.15 53.47
Llama3-8B-Instructlayout (Llama Team, 2024) 49.13 41.75 45.44 74.00 62.20 63.15 66.45
LayTextLLMzero (Lu et al., 2024) 65.50 37.40 51.45 72.00 45.50 82.00 66.50
Text + Layout + Image
LayoutLLMLlama2 (Luo et al., 2024) 74.25 55.73 64.99 78.65 62.21 70.97 70.61
DocLayLLMLlama2 (Ours) 72.83 55.04 63.94 78.67 70.81 83.27 77.58
DocLayLLMLlama3 (Ours) 78.40 58.55 68.48 84.12 71.34 84.36 79.94
Table 2: Comparison with OCR-dependent methods using metrics from LayoutLLM Luo et al. (2024). Avg. represents the average performance across different benchmarks, and - indicates the cleaned test set used in LayoutLLM. The suffix layout denotes the method from He et al. (2023), which provides LLM with layout features via OCR box coordinates in textual form. Bold indicates the best performance, and underline indicates the second-best one.
Method Document-oriented VQA KIE
@@
Llama2-7B-Chat (Touvron et al., 2023) 20.50 6.30 13.40 23.40 51.80 58.60 44.60
Llama3-8B-Instruct (Llama Team, 2024) 51.79 229.74 140.77 66.67 74.71 82.51 74.63
Text + Layout
Llama2-7B-Chatlayout (Touvron et al., 2023) 12.30 28.00 20.15 14.40 38.10 50.60 34.37
Llama3-8B-Instructlayout (Llama Team, 2024) 49.13 211.69 130.41 74.71 75.33 75.93 75.32
DocLLM (Wang et al., 2024) 69.50 264.10 166.80 51.80 67.60 91.90 70.43
LayTextLLMzero (Lu et al., 2024) 65.50 200.20 132.85 47.20 77.20 83.80 69.40
Text + Layout + Image
DocLayLLMLlama2 (Ours) 72.83 310.60 191.72 80.74 79.37 84.37 81.49
DocLayLLMLlama3 (Ours) 78.40 357.04 217.72 86.47 83.66 86.08 85.40
Table 3: Comparison with OCR-dependent methods using metrics from DocLLM Wang et al. (2024). indicates that the model uses the training set corresponding to the evaluation set during training.
4 Experiments
@@ 4.1 Datasets
During the pre-training stage, we conduct our pre-training tasks on several datasets, including RVL-CDIP (Pramanik et al., 2020), PubLayNet (Zhong et al., 2019), PubTabNet (Zhong et al., 2020), DocLayNet (Pfitzmann et al., 2022), DocBank (Li et al., 2020), DocILE (Simsa et al., 2023), and the Document Dense Description dataset from LayoutLLM (Luo et al., 2024). This results in a total of 3.1 million pairs of in-context learning data, which is significantly less than the volume used in previous OCR-dependent TDU methods. For the supervised fine-tuning stage, we apply CoT annealing on the layout-aware document QA data utilized in LayoutLLM (Luo et al., 2024).
We evaluate our DocLayLLM on widely used text-rich document benchmarks, including FUNSD (Jaume et al., 2019), CORD (Park et al., 2019), SROIE (Huang et al., 2019), POIE (Kuang et al., 2023), DocVQA (Mathew et al., 2021), InfoVQA (Mathew et al., 2022), VisualMRC (Tanaka et al., 2021), DeepForm (Svetlichnaya, 2020), KLC (Stanislawek et al., 2021), WTQ (Pasupat and Liang, 2015), and TabFact (Chen et al., 2020). To fairly compare DocLayLLM with other approaches, we follow the evaluation pipeline and metrics they used.
@@ 4.2 Implement Details
Our method is based on Llama3-8B-Instruct  [^1] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct (Llama Team, 2024). We also build a version using Llama2-7B-Chat  [^2] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf (Touvron et al., 2023) for comparison with other methods, denoted as DocLayLLMLlama3 and DocLayLLMLlama2 respectively. We initialize our visual encoder and layout encoder weights from LayoutLMv3 (Huang et al., 2022). During pre-training and fine-tuning, we use LoRA Hu et al. (2022) with a rank of 64. We train our model on 8 NVIDIA Tesla A100-80G GPUs, taking about 30 A100 days for pre-training and 6 A100 days for fine-tuning. For OCR results, we use the default OCR data from the official datasets or commercial OCR engines if not provided. Additional details can be found in Appendix A.1.
@@
LLaVAR (Zhang et al., 2023) 12.30 16.50 14.40 0.50 5.20 5.90 3.87
UniDoc (Feng et al., 2023b) 7.70 14.70 11.20 1.00 2.90 5.10 3.00
DocPedia (Feng et al., 2023a) 47.10 15.20 31.15 29.90 21.40 39.90 30.40
Monkey (Li et al., 2024b) 50.10 25.80 37.95 24.10 41.90 19.90 28.63
TextMonkey (Liu et al., 2024) 66.70 28.60 47.65 42.90 46.20 32.00 40.37
DocOwl 1.5 (Hu et al., 2024a) - - - - 48.30 51.80 -
TextSquare (Tang et al., 2024) - - - - 53.20 71.80 -
DocKylin (Zhang et al., 2024) 65.10 34.80 49.95 58.70 25.50 49.50 44.57
DocLayLLMLlama3 (Ours) 77.79 42.02 59.91 80.28 76.59 75.13 77.33
Table 4: Comparison with OCR-free methods using the precision metric from TextMonkey Liu et al. (2024). indicates that the model uses the training set corresponding to the evaluation set during training.
@@
DocPedia (Feng et al., 2023a) 47.10 15.20 - - - - - - - -
DocOwl (Ye et al., 2023a) 62.20 38.20 188.80 96.40 42.60 30.30 36.45 26.90 60.20 43.55
UReader (Ye et al., 2023b) 65.40 42.20 221.70 109.77 49.50 32.80 41.15 29.40 67.60 48.50
Monkey (Li et al., 2024b) 66.50 36.10 - - 40.60 32.80 - 25.30 - -
TextMonkey (Liu et al., 2024) 73.00 - - - 59.70 37.80 - 31.90 - -
DocKylin (Zhang et al., 2024) 77.30 46.60 319.90 147.93 64.20 - - 32.40 - -
CogAgent (Hong et al., 2024) 81.60 44.50 - - - - - - - -
DocOwl 1.5 (Hu et al., 2024a) 82.20 50.70 246.40 126.43 68.80 38.70 53.75 40.60 80.20 60.40
TextSquare (Tang et al., 2024) 84.30 51.50 - - - - 49.70 84.20 66.95
DocLayLLMLlama3 (Ours) 78.40 40.90 357.04 158.80 30.19 25.76 27.98 46.03 65.65 55.84
DocLayLLMLlama3 (Ours) 86.52 58.36 388.85 177.91 77.07 40.73 58.90 58.62 83.36 70.99
Table 5: Comparison with OCR-free methods using metrics from DocOwl 1.5 Ye et al. (2023a). indicates that the model uses the training set corresponding to the evaluation set during training.
@@ 4.3 Comparisons with SOTA OCR-dependent methods
We have compared our DocLayLLM with existing SOTA OCR-dependent methods. In order to exhibit the zero-shot inferring ability of OCR-dependent methods precisely, we chose their performances under the zero-shot scenario unless they are not provided. The experimental results are shown in Table 2 and Table 3.
It can be found that our method outperforms the baseline models, regardless of the type of baseline model used or whether the document’s layout information is provided to the baseline model. It validates the rationale behind our design of multi-modal extension.
Additionally, our method significantly outperforms previous SOTA methods, even with less pre-training data and fewer resources. As shown in Table 2, it surpasses the performance of LayoutLLM Luo et al. (2024) and LayTextLLM Lu et al. (2024) while using fewer computational resources by solely employing LoRA during training. Even with Llama2 as the backbone, our approach matches the performance of previous SOTA methods in the Document-oriented VQA task and significantly outperforms them in the KIE tasks, further demonstrating the efficiency and effectiveness of our approach.
Besides, our method exhibits robust zero-shot generalization capability without the need for fine-tuning on specific downstream tasks. From the comparison with DocLLM Wang et al. (2024) in Table 3, it is clear that our method, even without fine-tuning, significantly outperforms DocLLM across most datasets and achieves a considerable overall performance improvement. This validates our design, which preserves the strong generalization ability of the original LLM structure.
In summary, our DocLayLLM is an efficient and effective OCR-dependent TDU model, demonstrating powerful zero-shot generalization capability by encoding the OCR information of the document with LLM and setting the new SOTA performance compared to other OCR-dependent competitors.

### v3

@@ 3.2 CoT Pre-training
During the pre-training phase, we incorporate a wide range of tasks (as detailed in the appendix) to boost the LLM’s understanding of the visual and layout features inspired by Luo et al. [45]. Since the datasets related to these tasks are not designed explicitly for instruction-tuning, their annotations are not in a QA format. Therefore, we manually create QA-formatted instructions based on their annotations for training. However, simply converting task descriptions into $Question$ and annotations into $Answer$ is highly inefficient regarding data utilization. To address this, we incorporate CoT during the conversion to instruction-tuning data for specific tasks, asking LLM to solve our problem step-by-step thus enhancing the model’s comprehension abilities, as the examples illustrated in Fig 2. Our designed CoT primarily falls into the following three categories:
@@ 3.2.1 Layout-Type-Related CoT
We conduct a layout-type-related CoT for the layout analysis task. To determine the layout type of a given area, it is beneficial first to locate its region within the document and then infer based on the nearby layout element. When giving an area bounded by a bounding box, our layout-type-related CoT can be formulated as following steps:
@@
    Identify the text within the area and determine its approximate region within the document (e.g., top left, top right part of the document, etc).
@@ 3.2.2 Table-Structure-Aware CoT
Table-structure-aware CoT is primarily employed in the table analysis task. We have integrated the XY-Cut algorithm [48]—a widely-used technique for analyzing table structures. For instance, to locate an element at the $i$-th row and $j$-th column, the CoT process is outlined as follows:
@@
    Identify all the table headers and their respective locations. The header of the $j$-th column corresponds to the $j$-th element in the headers list.
@@
    Output all elements in the $j$-th column when told that the elements of the same column are roughly vertically aligned with their corresponding headers.
@@
Applying the above CoT, we break down a complex question into simple steps using the basic XY-Cut thought. This approach improves the LLM’s inference ability through this breakdown and enhances data utilization by effectively leveraging the overall structure of the table.
@@ 3.2.3 Geometry-Related CoT
This is used in geometric analysis to identify geometric relationships between two text segments. It assists LLM to inference with geometric relations. Specifically, when giving the text content of two lines:
@@
    Recover the bounding box of the given text segments.
@@
    Calculate the center point of each segment and determine their relative location by comparing their center points.
@@
This design advances the LLM’s arithmetic reasoning capabilities and helps it more accurately grasp the spatial relationships between different boxes, and also contributes to layout-type-aware CoT to find the nearest bounding boxes.
@@ 3.3 CoT Annealing
Data annealing involves gradually increasing the proportion of high-quality data towards the end of supervised fine-tuning to improve data utilization efficiency and augment the model’s reasoning abilities. This technique has been proven to be effective and is widely adopted in models like MiniCPM [19], DataComp-LM [31], and Llama-3.1 [42].
During the SFT stage, we employ data with a layout-aware CoT [45] that helps focus on potential regions of questions and answers within documents, thereby better leveraging the layout features learned during pre-training. It also allows the model to strengthen its reasoning capacity through step-by-step thinking. However, in the latter stages of SFT, our goal for LLM is to output precise and direct answers to questions. At this point, using CoT data may introduce unnecessary noise with a long response, and the overly uniform layout-aware CoT data might partially compromise the LLM’s generalization abilities.
To balance the strengths and weaknesses of such data, we propose CoT annealing, which reconsiders this data annealing issue from a CoT perspective. Specifically, we generate direct-answer data without CoT from the CoT data, treating it as high-quality data. During the SFT process, we gradually increase the proportion of this direct-answer data, thereby enhancing the LLM’s ability to respond with a clear and direct answer in the end. Initially, we only utilize the data with CoT, gradually adjusting the ratio of CoT to non-CoT data as the SFT process progresses. Ultimately, we conclude the training using only the data without CoT.
4 Experiments
Method Processed Data at Pre-training Document-oriented VQA KIE
@@ 4 Experiments
Llama2-7B-Chat [64] - 20.50 9.90 15.20 15.10 20.00 35.6 23.57
Llama3-8B-Instruct [42] - 51.79 47.77 49.78 68.56 52.31 61.24 60.70
Text + Box
Llama2-7B-Chatcoor [14] - 12.30 12.20 12.25 11.90 6.40 39.40 19.23
Llama3-8B-Instructcoor [14] - 49.13 41.75 45.44 74.00 62.20 63.15 66.45
LayTextLLMzero [43] 10.0M (323%) 65.50 37.40 51.45 72.00 45.50 82.00 66.50
Text + Box + Patch
LayoutLLMLlama2 [45] 5.7M (184%) 74.25 55.73 64.99 78.65 62.21 70.97 70.61
$DocLayLLM_{Llama2}$ (Ours) 3.1M (100%) 72.83 55.04 63.94 78.67 70.81 83.27 77.58
$DocLayLLM_{Llama3}$ (Ours) 3.1M (100%) 78.36 58.55 68.46 84.11 71.34 84.36 79.94
Table 1: Comparison with OCR-dependent methods using metrics from LayoutLLM [45]. Processed Data at Pre-training represents the total data volume processed at the pre-training stage, equivalent to dataset size multiplied by epoch count, with values in parentheses representing size ratios compared to DocLayLLM. Avg. represents the average performance across different benchmarks, and indicates the cleaned test set used in LayoutLLM. The suffix layout denotes using method from He et al. [14], which provides LLM with layout features via OCR box coordinates in textual form. Bold indicates the best performance, and underline indicates the second-best one.
Method Processed Data at Pre-training Document-oriented VQA KIE
DocVQA VisualMRC Avg. FUNSD CORD SROIE Avg.
Text
Llama2-7B-Chat [64] - 20.50 6.30 13.40 23.40 51.80 58.60 44.60
Llama3-8B-Instruct [42] - 51.79 229.74 140.77 66.67 74.71 82.51 74.63
Text + Box
Llama2-7B-Chatcoor [14] - 12.30 28.00 20.15 14.40 38.10 50.60 34.37
Llama3-8B-Instructcoor [14] - 49.13 211.69 130.41 74.71 75.33 75.93 75.32
LMDX [53] - - - - - 66.95 - -
DocLLM [67] 16.8M (542%) 69.50 264.10 166.80 51.80 67.60 91.90 70.43
LayTextLLMzero [43] 10.0M (323%) 65.50 200.20 132.85 47.20 77.20 83.80 69.40
Text + Box + Patch
$DocLayLLM_{Llama2}$ (Ours) 3.1M (100%) 72.83 310.60 191.72 80.74 79.37 84.37 81.49
$DocLayLLM_{Llama3}$ (Ours) 3.1M (100%) 78.35 357.04 217.70 86.47 83.66 86.08 85.40
Table 2: Comparison with OCR-dependent methods using metrics from DocLLM [67]. indicates that the model uses the training set of the benchmark during training.
@@ 4.1 Datasets
During the pre-training stage, we conduct our pre-training tasks on several datasets, including RVL-CDIP [13], PubLayNet [80], PubTabNet [81], DocLayNet [54], DocBank [32], DocILE [56], and the Document Dense Description data from LayoutLLM [45]. This results in a total of 3.1 million pairs of instruction tuning data, which is significantly less than the volume used in previous OCR-dependent TDU methods. For the supervised fine-tuning stage, we apply CoT annealing on the layout-aware document QA data utilized in LayoutLLM [45].
We evaluate our DocLayLLM on widely used text-rich document benchmarks, including FUNSD [22], CORD [50], SROIE [21], POIE [25], DocVQA [46], InfoVQA [47], VisualMRC [59], DeepForm [58], KLC [57], WTQ [51], and TabFact [4]. For a fair comparison with other approaches, we use the default OCR data from the official datasets otherwise commercial OCR engines  [^1] https://www.textin.com/ if not provided, and follow their evaluation pipeline and metrics.
@@ 4.2 Implement Details
Our method is based on Llama3-8B-Instruct [42], with a Llama2-7B-Chat [64] version for better comparison with other methods. They are denoted as DocLayLLMLlama3 and DocLayLLMLlama2 respectively. We initialize $VE$ and $LE$ with weights from LayoutLMv3 [20] to leverage document knowledge acquired during its pre-training, while $VP$ and $LP$ are randomly initialized. Besides, the image resolution is fixed at 224*224, resulting in 196 visual patch tokens, to capture an overview of document layout while bringing less computation burden. During pre-training and fine-tuning, we use LoRA [18] with a rank of 64 on LLM and keep $VE$, $VP$, $LE$, and $LP$ trainable, leading to few tuning parameters. We train our model on 8 NVIDIA Tesla A100-80G GPUs, taking about 30 A100 days for pre-training and 6 A100 days for supervised fine-tuning. Additional details can be found in the appendix.
@@
LLaVAR [79] 12.30 16.50 14.40 0.50 5.20 5.90 3.87
UniDoc [10] 7.70 14.70 11.20 1.00 2.90 5.10 3.00
DocPedia [9] 47.10 15.20 31.15 29.90 21.40 39.90 30.40
Monkey [38] 50.10 25.80 37.95 24.10 41.90 19.90 28.63
TextMonkey [41] 66.70 28.60 47.65 42.90 46.20 32.00 40.37
DocOwl 1.5 [17] - - - - 48.30 51.80 -
TextSquare [61] - - - - 53.20 71.80 -
DocKylin [78] 65.10 34.80 49.95 58.70 25.50 49.50 44.57
$DocLayLLM_{Llama3}$ (Ours) 77.79 42.02 59.91 80.28 76.59 75.13 77.33
Table 3: Comparison with OCR-free methods using the precision metric from TextMonkey [41]. indicates that the model uses the training set of the benchmark during training.
@@
DocPedia [9] 47.10 15.20 - - - - - - - -
DocOwl [75] 62.20 38.20 188.80 96.40 42.60 30.30 36.45 26.90 60.20 43.55
UReader [76] 65.40 42.20 221.70 109.77 49.50 32.80 41.15 29.40 67.60 48.50
Monkey [38] 66.50 36.10 - - 40.60 32.80 36.70 25.30 - -
TextMonkey [41] 73.00 - - - 59.70 37.80 48.75 31.90 - -
DocKylin [78] 77.30 46.60 319.90 147.93 64.20 - - 32.40 - -
CogAgent [16] 81.60 44.50 - - - - - - - -
DocOwl 1.5 [17] 82.20 50.70 246.40 126.43 68.80 38.70 53.75 40.60 80.20 60.40
TextSquare [61] 84.30 51.50 - - - - - 49.70 84.20 66.95
$DocLayLLM_{Llama3}^{\ddagger}$ (Ours) 86.52 58.36 327.91 157.60 77.07 40.73 58.90 58.62 83.36 70.99
Table 4: Comparison with OCR-free methods using the evaluation framework and metrics from DocOwl 1.5 [17]. All methods, including our $DocLayLLM_{Llama3}^{\ddagger}$ variant, are exposed to the train set corresponding to the evaluation set during training.
@@ 4.3 Comparisons with SOTA OCR-dependent methods
We have compared our DocLayLLM with existing SOTA OCR-dependent methods. In order to exhibit the zero-shot inferring ability precisely, we chose their performances under the zero-shot scenario unless they are not provided.
The experimental results are shown in Table 1 and Table 2. Our method outperforms the baseline models, regardless of their type or whether the document’s layout information is provided to them. It validates the rationale behind our design of multi-modal extension.
Additionally, our method significantly outperforms previous SOTA methods despite fewer training resources. As shown in Table 1, it surpasses the performance of LayoutLLM [45] and LayTextLLM [43] though requiring fewer computational resources by using less data and tuning few parameters at training. Even with Llama2 as the backbone, our approach matches the performance of previous SOTA methods in the Document-oriented VQA task and significantly outperforms them in the KIE tasks, further demonstrating the efficiency and effectiveness of our approach.
Besides, our method exhibits robust zero-shot generalization capability without fine-tuning with specific downstream tasks. From the comparison with DocLLM [67] in Table 2, it is clear that our method, even without fine-tuning, significantly outperforms DocLLM across most datasets and achieves a considerable overall performance improvement. This validates our design, which has preserved the strong generalization ability of the original LLM structure.
In summary, DocLayLLM serves as an efficient extension of LLMs. It demonstrates strong zero-shot generalization and effective document understanding by encoding documents within a unified model. Even with limited training resources, DocLayLLM achieves new SOTA performance compared to OCR-dependent competitors.